OpenStudy (anonymous):

How do you solve this indefinite integral... (sen^5x/sqrt(cosx)) dx

OpenStudy (anonymous):

\[\int\limits \frac{ sen ^{5}x }{ \sqrt{cosx} }\]

OpenStudy (sirm3d):

\[u=\cos x\] rewrite \[\sin^5 x=(sin^2 x) (\sin^2 x) (\sin x)\] use a trigonometric identity on \[\sin^2 x\]

OpenStudy (anonymous):

¿?